OpenStudy (anonymous):

Multiply sqrt[3]{9x^5}*sqrt[3]{24x} Simplify the product

OpenStudy (anonymous):

\[\sqrt[3]{9x^5}*\sqrt[3]{24x}\]

Parth (parthkohli):

\( \color{Black}{\Rightarrow \sqrt[3]{x} \times \sqrt[3]{y} = \sqrt[3]{xy} }\) Multiply both and put it under the radical.

Parth (parthkohli):

\( \color{Black}{\Rightarrow \large \sqrt[3]{9x^5 \times 24x} }\)

OpenStudy (lgbasallote):

you have same indeces sp it's just \[\sqrt[3]{(9x^5) \times (24x)}\]

OpenStudy (inkyvoyd):

btw, for texing, put it in \(

OpenStudy (inkyvoyd):

and \)

OpenStudy (lgbasallote):

so it's equal to \[\sqrt[3]{216x^6}\] got it?

OpenStudy (lgbasallote):

i just multiplied 9x^5 and 24x

Parth (parthkohli):

It can be simplified. \( \color{Black}{\Rightarrow 6x^2}\)

OpenStudy (lgbasallote):

216 and x^6 are perfect cubes (i believe)

OpenStudy (inkyvoyd):

Guys, it is a very bad idea to do that.

Parth (parthkohli):

\( \color{Black}{\Rightarrow \sqrt[3]{216} \times \sqrt[3]{x^6} = 6 \times x^2 = 6x^2 }\)

OpenStudy (anonymous):

Why is it a bad idea?

OpenStudy (lgbasallote):

coz he's a troll...the blank pic is a giveaway isnt it ;D

OpenStudy (anonymous):

@lgbasallote haven't i seen you somewhere before?

Parth (parthkohli):

lol he's a troll he put myin's pic

OpenStudy (inkyvoyd):

\(\sqrt[3]{9x^5}*\sqrt[3]{24x}\) =\(\sqrt[3]{x^5*x}*\sqrt[3]{24*9}\) =\(\sqrt[3]{x^5*x}*\sqrt[3]{2^2*3*3^2}\) =\(\sqrt[3]{x^5*x}*\sqrt[3]{2^2*3^3}\) =\(\sqrt[3]{x^5*x}*2*3\) =\(6\sqrt[3]{x^5*x}\) =\(6\sqrt[3]{x^6}\) =\(6x^2\)

OpenStudy (lgbasallote):

hmm no @satellite73 i don't believe we've met

OpenStudy (inkyvoyd):

Never multiply something that you will factor. You are wasting time by multipling it and still having to factor it. it is ALWAYS a better idea to factor something, so taking radicals of it is easier.

OpenStudy (inkyvoyd):

@cubanito305

OpenStudy (anonymous):

is there a tutoral i can see on factoring, i always have trouble with that >.<

OpenStudy (inkyvoyd):

Well, I usually check if itis divisible by 2 or 3. Remember, if the sum of the digits is a multiple of 3, it is divisible by 3. If it is even, it is divisible by 2.

OpenStudy (lgbasallote):

my factoring tutorial is on qadratic expressions

OpenStudy (inkyvoyd):

if it ends with a 5 or 0, it is divisible by 5. Divisbility by 7 is checkable too, but I usually do the whole long division.

OpenStudy (inkyvoyd):

factor any number out of the expression if possible, even if you know that you can't simplify it.

OpenStudy (anonymous):

thanks